3 ms·
Intuitively it might make sense to add the probabilities, but it doesn't actually work out that way[0]. The parent used the law of total probability to solve fo
by elsherbini 9y ago
Intuitively it might make sense to add the probabilities, but it doesn't actually work out that way[0]. The parent used the law of total probability to solve for the probability of hitting at least once because it's simpler:
P(hits >= 1) = 1 - P(hits = 0).
Solving for the left side of the equation directly gets trickier:
P(hits >= 1) = P(hits = 1) + P(hits = 2) + ... + P(hits = k)
P(hits = 1) = 100 chose 1 * 0.01^1 * 0.99^99
...
p(hits = n) = 100 chose n * 0.01^n *0.99^(100-n)
Taking that sum where n=100 yields the same as
1 - P(hits = 0)
= 1 - (100 chose 0 *0.1^0 *0.99^100)
= 1 - 0.99^100
[0] https://en.wikipedia.org/wiki/Binomial_distribution https://en.wikipedia.org/wiki/Binomial_distribution